OpenStudy (anonymous):

Find the length of the radius of the following circle. (x + 2)² + y² = 10 √(10) 10

OpenStudy (anonymous):

is it 10?

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (xapproachesinfinity):

try to write 10 as something squared like r^2

OpenStudy (xapproachesinfinity):

if you looked at the previous post i said the general form is (x-a)^2+(y-b)^2=r^2 this means r^2=10 in your case

OpenStudy (xapproachesinfinity):

solve r^2=10

OpenStudy (anonymous):

5?

OpenStudy (xapproachesinfinity):

no how do you solve x^2=4

OpenStudy (anonymous):

thnx

OpenStudy (xapproachesinfinity):

did you find the answer to r^2=10?